OpenStudy (anonymous):

Describe the transformations required to obtain the graph of the function f(x) from the graph of the function g(x). f(x) = 9 cos x ; g(x) = cos x

OpenStudy (cwrw238):

1/9 f(x) = g(x) thats a stretch factor (1 / 1/9) = 9 parallel to the y-axis

OpenStudy (anonymous):

f(x) is nine times bigger than g(x), so f(x)'s line is 9 times higher than g(x).

OpenStudy (cwrw238):

thats correct

OpenStudy (cwrw238):